OpenStudy (anonymous):

A brick is thrown upward from the top of a building at an angle of 25 degrees to the horizontal with an initial speed of 15 m/s. If the brick is in flight for 3 s, how tall is the building? Answer is 25 m. How?

OpenStudy (anonymous):

\[V _{x}=15\cos25=13.59m/s\] \[V _{y}=15\sin25=6.34m/s\] \[\Delta y=v _{0}t+\frac{ 1 }{ 2 }at ^{2}\] \[\Delta y=(6.34)(3)+ (4.9)(3) ^{2}\]=63.12 What am I doing wrong?

OpenStudy (anonymous):

check 4th line....

OpenStudy (anonymous):

of ur solution...

OpenStudy (anonymous):

your problem ...whether acceleration due to gravity will be + or -? check this link...scroll down in this page...U'll see something about frame of reference.

OpenStudy (anonymous):